OpenStudy (anonymous):

6. The equation of line AB is y = 5x + 1. Write an equation of a line parallel to line AB in slope-intercept form that contains point (4, 5). (1 point

OpenStudy (texaschic101):

y = 5x + 1 the slope here is 5 a parallel line will have the same slope now we will use y = mx + b (4,5)...x = 4 and y = 5 slope(m) = 5 now we sub, and since we already know the slope, we will find the y intercept, which is b. 5 = 5(4) + b 5 = 20 + b 5 - 20 = b 15 = b so our parallel equation is : y = 5x + 15 any questions ?

OpenStudy (anonymous):

22. A cone has a volume of 5 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of? (1 point) 15 in 3 20 in 3 25 in 3 30 in 3 23. Find the diameter of a cone that has a volume of 104.67 cubic inches and a height of 4 inches. Use 3.14 for pi. (1 point) 4 inches 6 inches 8 inches 10 inche